Please help me solve this linear equation:
(x-9)^2+y^2=25
4x-3y=61

Answer by Alan3354(69443)

Please help me solve this linear equation:
(x-9)^2+y^2=25
4x-3y=61 --> 3y = 4x-61
y = (4x-61)/3
--------------
Sub for y
(x-9)^2+((4x-61)/3)^2=25
Multiply by 9
9*(x-9)^2 + (4x-61)^2 = 225
9x^2 - 162x + 729 + 16x^2 - 488x + 3721 = 225
25x^2 - 650x + 4450 = 225
divide by 25
x^2 - 26x + 178 = 9
x^2 - 26x + 169 = 0
(x-13)^2 = 0
x = 13, x = 13
Why do you call that linear? It's a quadratic.
4*13 - 3y = 61
y = -3